OpenStudy (anonymous):

Solve: |2n - 3| = 5 A. {-1, 4} B. {1, 4} C. {-1, 1} I pick b but i want to make sure D. {1, -4}

OpenStudy (anonymous):

I would go with A |2(-1) - 3| = 5 |-2 - 3| = 5 |-5| = 5 and |2(4) - 3| = 5 |8 - 3| = 5 |5| = 5